McCormick & Company is considering a project that requires an initial investment of $24 million to build a new plant and purchase equipment. The investment will be depreciated as a modified accelerated cost recovery system (MACRS) seven-year class asset. The new plant will be built on some of the company's land, which has a current, after-tax market value of $4.3 million. The company will produce bulk units at a cost of $130 each and will sell them for $420 each. There are annual fixed costs of $500,000. Unit sales are expected to be $150,000 each year for the next six years, at which time the project will be abandoned. At that time, the plant and equipment is expected to be worth $8 million (before tax) and the land is expected to be worth $5.4 million (after tax). To supplement the production process, the company will need to purchase $1 million worth of inventory. That inventory will be depleted during the final year of the project. The company has $100 million of debt outstanding with a yield to maturity of 8 percent, and has $150 million of equity outstanding with a beta of 0.9. The expected market return is 13 percent, and the risk-free rate is 5 percent. The company's marginal tax rate is 40 percent. Should the project be accepted? questions: 6. Create an after-tax cash flow timeline. (what's the formula?) 7. What are the total expected cash flows at the end of year six? The $4.3 million is an opportunity cost and must be included at date zero as a cash outflow. If the project is accepted, however, the land can be sold in six years for $5.4 million. 8. Find the NPV using the after-tax WACC as the discount rate. 9. Find the IRR. 10. Should the project be accepted? Discuss whether NPV or IRR creates the best decision rule.
Answer #1

net cash from sale of land in year 6 = $5.4 million - ($5.4 million - $4.3 million) * (40%) = $4,960,000

accumulated depreciation on plant upto end of year 6 = $20,786,400

book value of plant at end of year 6 = $24 million - $20,786,400 = $3,213,600

sale price of plant at end of year 6 = $8 million

net cash from sale of plant = $8 million - ($8 million - $3,213,600)*(40%) = $6,085,440

net cash flow in each year = income after taxes + depreciation

f J8-G8+K8+L8+M8 N8 A B G H M N C Net Cash from Sale of plant and Equipment Flow Investment PV of Net Cash Flow Net Cash Inve

Present value of each net cash flow, discounted at the after-tax WACC discount rate, is the PV of each net cash flow. The sum of the PVs of each net cash flow is the NPV of the project

WACC = (cost of debt * weight of debt) + (cost of equity * weight of equity)

cost of debt = after tax cost of debt = YTM * (1 - tax rate) = 8% * (1 - 40%) = 4.8%

cost of equity = risk free rate + (beta * market risk premium) = 5% + (0.9 * (13% - 5%)) = 12.2%

weight of debt = $100 million / ($100 million + $150 million) = 0.40

weight of equity = $150 million / ($100 million + $150 million) = 0.60

WACC = (4.8% * 0.40) + (12.2% * 0.60) = 9.24%

NPV of the project is $99,112,261

IRR of the project is 76.22%

The project should be accepted as the NPV is positive, and the IRR is higher than the WACC

f 07 N7 / 1.0924 A7 G K N Net Cash from Sale of Investment Net Cash from Sale of land Net Cash Investment plant and land Inco